# Sign-In to the Lottery: Reparameterizing Sparse Training From Scratch

### v1

@@ Abstract
The performance gap between training sparse neural networks from scratch (PaI) and dense-to-sparse training presents a major roadblock for efficient deep learning. According to the Lottery Ticket Hypothesis, PaI hinges on finding a problem specific parameter initialization. As we show, to this end, determining correct parameter signs is sufficient. Yet, they remain elusive to PaI. To address this issue, we propose Sign-In, which employs a dynamic reparameterization that provably induces sign flips. Such sign flips are complementary to the ones that dense-to-sparse training can accomplish, rendering Sign-In as an orthogonal method. While our experiments and theory suggest performance improvements of PaI, they also carve out the main open challenge to close the gap between PaI and dense-to-sparse training.
Sparse Training, Pruning at Initialization, Lottery Tickets, Sign Flips, Deep Learning Efficiency, Overparameterization, Continuous Sparsification
@@ 1 Introduction
As neural network sizes scale to billions of parameters (Hoffmann et al., 2022; Kaplan et al., 2020), building efficient training pipelines is of paramount importance. Neural network sparsification plays a pivotal role in enabling such efficient pipelines (Frantar & Alistarh, 2023; Jaiswal et al., 2024). State-of-the-art sparsification methods rely on eliminating parameters during training (Peste et al., 2021; Kuznedelev et al., 2024; Kusupati et al., 2020) or by exploring multiple sparse masks dynamically (Evci et al., 2020; Lasby et al., 2023; Chen et al., 2021) to find high-performing sparse networks. The success of these methods can be attributed to finding an optimal combination of parameters and sparse masks (Paul et al., 2023). However, identifying the optimal combination requires training the dense network before it can be sparsified such that information from the dense network can be inherited by the subsequent sparse network (Gadhikar & Burkholz, 2024a).
@@
It is then no surprise that the most successful sparsification methods rely on dense training phases. Dense-to-sparse methods like gradual pruning (Han et al., 2015) or continuous sparsification (Kusupati et al., 2020) explicitly start from a dense network. Dynamic sparse training methods like RiGL (Evci et al., 2020) utilize dense gradient information. However, this comes at the cost of high computational and memory demands during the dense training phase. Moreover, the benefits of sparsification become mostly available towards the end of training or during inference.
@@
Is there a way to train sparse networks from the start? The Lottery Ticket (LT) Hypothesis (Frankle & Carbin, 2019) posits that there exist sparse networks at random initialization which enjoy optimal parameter-mask coupling and can achieve the same performance as their dense counterpart. Finding such sparse networks at initialization would unlock the benefits of sparse training from scratch. A plethora of pruning at initialization (PaI) methods have attempted to find such sparse networks at initialization (Lee et al., 2019; Tanaka et al., 2020; Wang et al., 2020; Pham et al., 2023; Liu et al., 2021), however, they struggle to match the generalization performance of dense-to-sparse training methods.
@@
Understanding and overcoming the limitations of PaI remains an open research problem and is the focus of our work. The main challenge for PaI methods is to find highly performant parameter-mask combinations by leveraging task-specific information efficiently (Kumar et al., 2024; Frankle et al., 2021). The problem is thus comprised of two parts: a) learning a mask and b) optimizing the weights corresponding to the mask, which is hard if the mask is extremely sparse. Here we are less concered about learning a mask, but instead focus on the hard problem of optimizing parameter-mask combinations, given a mask at initialization. At the center of our investigation is the question:
@@
What hinders training a given sparse network and can the gap between dense-to-sparse training and PaI be closed?
@@
The key to addressing this question lies in promoting parameter sign flips. Echoing the analysis of learning rate rewinding (LRR) by Gadhikar & Burkholz (2024b), through a broader empirical investigation of state-of-the-art dense-to-sparse training methods, we identify a critical phenomenon responsible for finding optimal parameter-mask combinations: the ability to flip parameter signs during training. We term this phenomenon sign alignment, as the signs encode all relevant information additionally to the mask (and data) to enable effective training from scratch.
Moreover, as we show, sign alignment occurs early in the dense-to-sparse training process, promoting convergence toward flatter minima. In contrast, PaI faces challenges in stabilizing sign flips. In line with this observation, it was shown in a theoretically tractable setting of a two-layer network consisting of one neuron that training a sparse network from scratch fails to flip signs from an unfavorable initialization (Gadhikar & Burkholz, 2024b).
@@
Motivated by this insight, we propose a way to flip and learn correct signs during sparse training from scratch in a higher number of cases (see Figure 2). We achieve this via Sign-In, a reparameterization $m\odot w$ of the non-masked parameters, providing them with an inner degree of freedom (where $m$ and $w$ are continuous parameters). Note that this reparameterization, depending on the inner scaling, has a bias towards sparsity. This has been used by the continuous sparsification methods Spred (Ziyin & Wang, 2023) and PILoT (Jacobs & Burkholz, 2025). In contrast, we do not use the reparameterization for sparsification but for improving the optimization geometry to enable sign alignment and thus seek to counteract its sparsity bias. We propose an inner scaling that promotes sign flips by rebalancing network layers and, correspondingly, the speed in which they learn.
@@
This provably recovers a different sign case than dense training and could thus be seen as complementary (see Figure 2). Moreover, in combination with a moderate degree of over-parameterization, which tends to be common in PaI, Sign-In fully solves the one-hidden-neuron problem. In line with this insight, we verify in experiments that existing dense-to-sparse methods can be improved by utilizing Sign-In. Most importantly, we demonstrate across multiple experiments that Sign-In significantly enhances the performance of sparse training from scratch, including those using random masks. We attribute this improvement to better sign alignment and the ability to discover flatter minima. However, as we show, Sign-In cannot replicate the same sign flipping mechanism as that of initial dense training, and is still not competitive with the accuracy of dense-to-sparse methods.
@@
Contributions
1) We empirically show that dense-to-sparse training methods identify trainable parameter-mask combinations via early sign alignment allowing them to find flatter minima and achieve state-of-the-art performance. 2) We provide evidence for the following hypothesis: Across dense-to-sparse training masks, the learned signs define an initialization that enables competitive sparse training from scratch. A fully random initialization (that is decoupled from the mask) performs on par or worse than PaI (including random pruning). 3) To learn signs more effectively, we propose Sign-In, a weight reparameterization, which can provably recover correct signs in a case that is complementary to overparameterization. 4) Extensive empirical evaluation shows that Sign-In can significantly improve the performance of sparse training from scratch by improving sign alignment and finding flatter minima.
@@ 2 Related Work
Training sparse neural networks from scratch is a hard problem (Kumar et al., 2024), but Frankle & Carbin (2019) posit that there exist initializations that make sparse networks trainable, i.e., Lottery Tickets. They occur as subnetworks of random source networks. However, their approach to find such initializations by iterative magnitude pruning (IMP) is computationally expensive and its performance does not carry over well to larger models, even though lottery tickets should exist in theory (da Cunha et al., 2022; Burkholz, 2022a, b; Burkholz et al., 2022; Burkholz, 2024). Renda et al. (2020) improved its accuracy by Weight Rewinding (WR) and Learning Rate Rewinding (LRR). Both WR and LRR are similarly expensive, but inherit the initialization from denser networks.
Dense-to-sparse methods
@@
While dense training has been conjectured to be a critical factor to enter a good loss basin Paul et al. (2023) or quadrant (Gadhikar & Burkholz, 2024a), other methods aim to reduce its duration. Gradual pruning methods like AC/DC (Peste et al., 2021), CAP (Kuznedelev et al., 2024), and WoodFisher (Singh & Alistarh, 2020) prune based on a parameter importance score, interspersed with training. Dynamic sparse training methods like RiGL (Evci et al., 2020; Lasby et al., 2023; Chen et al., 2021) dynamically adapt a sparse mask during training using dense gradients to inform reactivation of weights. Continuous sparsification methods such as STR (Kusupati et al., 2020), spred (Ziyin & Wang, 2023), and PILoT (Jacobs & Burkholz, 2025) reparameterize a dense neural network to induce sparsity. Although such dense training phases seem to be critical to learn good parameter signs (Gadhikar & Burkholz, 2024b), we propose a complementary method, which also works for pruning at initialization.
Pruning at initialization (PaI)
@@
A variety of methods have been proposed to sparsify networks at initialization based on different scores that assess parameter importance like saliency in SNIP (Lee et al., 2019) and GraSP (Wang et al., 2020), signal flow in Synflow (Tanaka et al., 2020) and maximal paths (Pham et al., 2023). However, randomly pruned sparse network can perform at par with these methods (Liu et al., 2021; Gadhikar et al., 2023; Ma et al., 2021; Su et al., 2020), calling our ability to identify problem specific masks into question (Frankle et al., 2021; Jain et al., 2024). Kumar et al. (2024) characterize the challenge from an information-theoretic lens, showing that initial dense training, in contrast to PaI, increases effective model capacity. Our goal is not to address the mask identification problem but to improve sparse training given a mask, by a reparameterization.
Reparameterization
@@
Reparameterizations have been studied from the perspective of implicit bias (Chizat & Bach, 2020; Li et al., 2022; Woodworth et al., 2020; Gunasekar et al., 2020, 2017; Chou et al., 2024; Vaškevičius et al., 2019; Li et al., 2023; Zhao et al., 2022; Li et al., 2021; Jacobs & Burkholz, 2025). It explains how overparameterization can improve generalization and how this is affected by various training factors such as the initialization, learning rate, noise, and explicit regularization. An important assumption in these works is that the activation function is linear. This simplification enables the reparameterization to be expressed using mirror flow or Riemannian gradient flow frameworks (Li et al., 2022). In contrast, we employ a reparameterization specifically to induce sign flips in the presence of a non-linear activation function. Our proposed reparameterization has been applied to induce continuous sparsification (Ziyin & Wang, 2023; Jacobs & Burkholz, 2025) with the help of explicit regularization. To combat this trend, we rescale the parameters dynamically exploiting a rescale invariance.
@@ 3 Correct Signs Facilitate Sparse Training
(a) Dense-to-sparse
(b) Sparse (PaI)
Figure 1: Sign learning in dense-to-sparse methods compared with training from scratch for a $90\%$ sparse ResNet50 on ImageNet. (a) The majority of signs are flipped early, upto warmup. Moreover, signs stabilize after warmp-up. (b) In contrast, signs do not stabilize after warmup for sparse training from scratch with different masks.
Paul et al. (2023) have established that iterative pruning methods like Learning Rate Rewinding (LRR) and Weight Rewinding (WR) rely on the effective exploration of the loss landscape during dense overparameterized training. Dense-to-sparse training inherits its performance by training overparameterized in the beginning, identifying a good loss basin. Gadhikar & Burkholz (2024a) have shown that this loss basin can be characterized by the weight signs and thus the quadrant of the solution. Our first objective is to show that this finding is not unique to LRR. In fact, it is a common phenomenon that points to a critical limitation of PaI to learn signs.
Our experiments reveal two key findings: (i) Initializing a network with an aligned sign and mask is sufficient for successful sparse training. (ii) Sign alignment happens early during training for dense-to-sparse methods and guides the training trajectory towards flatter minima. Specifically, we study three well-established sparsification methods to draw insights on sign alignment: magnitude based pruning with AC/DC (Peste et al., 2021), continuous sparsification with STR (Kusupati et al., 2020), and dynamic sparse training with RiGL (Evci et al., 2020). These methods leverage information from dense training by alternating dense and sparse training, gradual sparsification by soft thresholding, or by retaining dense gradient information, respectively. For comparison with sparse training, we train a randomly pruned sparse mask from scratch, which has been shown to be competitive with (and even better than) other PaI methods (Liu et al., 2021) and serves as a strong baseline. Noteworthy is the fact that the random mask contains no information about the task. Even though other masks should therefore have an advantage, as we will see later, they cannot shine through because of limited trainability.
@@
Method AC/DC RiGL STR
Baseline $\underline{74.68}$ $\underline{73.75}$ $\underline{73.65}$
Epoch $10$ sign + learned mask $73.97$ $73.32$ $71.7$
Epoch $30$ sign + learned mask $\mathbf{74.89}$ $73.7$ $\mathbf{73.91}$
Final sign + learned mask $74.88$ $\mathbf{73.74}$ $73.77$
Epoch $10$ mag + learned mask $70.93$ $72.30$ $68.01$
Final mag + learned mask $70.94$ $72.40$ $68.35$
Random init + learned mask $70.6$ $71.89$ $68.38$
@@
Signs are learned early To study the alignment between the sparse mask and parameter signs, we track them during dense-to-sparse training of a ResNet50 on ImageNet (see Section 6 for details). As shown in Figure 1(a), the initial signs and final learned signs are significantly different since approximately $50\%$ of the signs are flipped for all methods (blue bar), i.e., the signs at initialization have a close to random overlap with the final learned signs. As training progresses, upto the warmup phase ($10$-th epoch), a significant number of signs flip, as shown by the red bar in Figure 1(a). After the warmup phase, fewer signs flip in addition (green bar in Figure 1(a)). In comparison, training the same masks from scratch starting from a random re-initialization (which is decoupled from the dense-to-sparse training procedure) does not identify stable signs early. The performance of these masks is at par with a random mask trained from scratch (see Table 10 in the appendix) in the absence of sign alignment. Moreover, the sign flip statistics show a similar pattern as training a random mask from scratch. A significant number of sign flips occur after warmup, and the overlap between initialization, warmup and the final signs is closer to random (see Figure 1(b)). Across all three dense-to-sparse baselines, we observe that signs are learned early during training and remain largely stable after warmup. In contrast, sparse training does not find stable signs to enable sign alignment, reflecting in its worse performance.
@@
Signs are more informative than magnitude As we verify next, the critical information about the task that is missing from the mask are the parameter signs. To do so, we use signs learned at different stages during training, the learned mask and random magnitudes for each method to initialize a sparse network. As reported in Table 1, final signs combined with the mask are sufficient for sparse training and achieve performance comparable to the baseline. Moreover, since the signs are identified as early as the $10$-th epoch during training, these early signs (combined with the mask) recover baseline performance with an error of $\approx 1\%$. Full baseline performance can be recovered by using signs from the $30$-th epoch. In contrast, if we initialize the sparse network with magnitudes learned at intermediate epochs, learned mask and random signs, we barely improve over training a randomly intialized sparse mask from scratch Table 1.
@@
Our results not only suggest that sign alignment is sufficient for sparse training, it also happens early during dense-to-sparse training. Sparse training from scratch with a random initialization falls short in comparison.
@@
Flatter loss landscapes with dense-to-sparse methods To dive deeper into the benefits of sign alignment for trainability, we track the maximum eigenvalue of the Hessian of the loss during training as a proxy of sharpness (Jiang et al., 2019). It can also be interpreted as a proxy measure for the tractability (or conditioning) of the optimization problem. As shown in Figure 5(b), the dense-to-sparse baselines AC/DC, STR and RiGL have smaller sharpness indicating a flatter loss landscape once signs align during training. In comparison, the random sparse network converges to a sharper minimum.
@@
Our observations lead to the conclusion that dense-to-sparse methods are effective because they inherit critical information from dense training in the form of parameter signs during the early training phase. These signs play a crucial role in dictating the training trajectory of the sparse network. In contrast, this mechanism is absent in sparse training, resulting in poor performance when using a randomly pruned mask and a sharper loss landscape during training, which indicates convergence to an inferior loss basin.

### v2

# Sign-In to the Lottery: Reparameterized Sparse Training

@@ Abstract
The performance gap between training sparse neural networks from scratch (PaI) and dense-to-sparse training, presents a major roadblock for efficient deep learning. According to the Lottery Ticket Hypothesis, PaI hinges on finding a problem specific parameter initialization, given a sparse mask. As we show, to this end, determining correct parameter signs is sufficient. Yet, they remain elusive to PaI. To address this issue, we propose Sign-In, which employs a dynamic reparameterization that provably induces sign flips. Such sign flips are complementary to the ones that dense-to-sparse training can accomplish, rendering Sign-In as an orthogonal method. While our experiments and theory suggest performance improvements of PaI, they also carve out the main open challenge to close the gap between PaI and dense-to-sparse training.
@@ 1 Introduction
As neural network sizes scale to billions of parameters, building efficient training pipelines is of paramount importance. Sparsification plays a pivotal role in enabling such efficient pipelines (Frantar & Alistarh, 2023). State-of-the-art sparsification methods rely on eliminating parameters during training (Kuznedelev et al., 2024). The success of these methods can be attributed to finding an optimal combination of parameters and sparse masks (Paul et al., 2023). However, identifying them requires training a dense network before it can be sparsified such that information from the dense network can be inherited by the subsequent sparse network (Gadhikar & Burkholz, 2024).
@@
It is then no surprise that the most successful sparsification methods rely on dense training phases. Dense-to-sparse methods like gradual pruning (Han et al., 2015) or continuous sparsification (Kusupati et al., 2020) explicitly start from a dense network. However, this comes at the cost of high computational and memory demands during the dense training phase. Moreover, the benefits of sparsification become mostly available towards the end of training or during inference.
@@
Is there a way to train sparse networks from the start, given a sparse mask? The Lottery Ticket (LT) Hypothesis (Frankle & Carbin, 2019) posits that there exist sparse networks at random initialization which enjoy optimal parameter-mask coupling and can achieve the same performance as their dense counterpart. Finding such sparse networks at initialization would unlock the benefits of sparse training from scratch. A plethora of pruning at initialization (PaI) methods have attempted to find such sparse networks at initialization (Lee et al., 2019; Tanaka et al., 2020; Wang et al., 2020), however, they struggle to match the generalization performance of dense-to-sparse training methods.
@@
Understanding and overcoming the limitations of PaI remains an open research problem and is the focus of our work. The main challenge for PaI methods is to find highly performant parameter-mask combinations by leveraging task-specific information efficiently (Kumar et al., 2024; Frankle et al., 2021). The problem is thus comprised of two parts: a) learning a mask and b) optimizing the weights corresponding to the mask, which is hard if the mask is extremely sparse. Here we are less concerned about learning a mask, but instead focus on the hard problem of optimizing parameter-mask combinations, given a mask at initialization. At the center of our investigation is the question:
@@
What hinders training a sparse mask, and can the gap between PaI and dense-to-sparse be closed?
@@
The key to addressing this question lies in promoting parameter sign flips. Echoing the analysis of learning rate rewinding (LRR) by Gadhikar & Burkholz (2024), through a broader empirical investigation of dense-to-sparse training methods, we identify a critical phenomenon responsible for finding optimal parameter-mask combinations: the ability to flip parameter signs during training. We term this phenomenon sign alignment, as the signs encode all relevant information additionally to the mask (and data) to enable effective training from scratch. Moreover, as we show, sign alignment occurs early in the dense-to-sparse training process, promoting convergence toward flatter minima. In contrast, PaI faces challenges in stabilizing sign flips. In line with this observation, it was shown in a theoretically tractable setting of one-hidden-neuron that training a sparse network from scratch fails to flip signs from a bad initialization (Gadhikar & Burkholz, 2024).
@@
Motivated by this insight, we propose a way to flip and learn correct signs during sparse training from scratch in a higher number of cases (see Figure˜2). We achieve this via Sign-In, a reparameterization $m\odot w$ of the non-masked parameters, providing them with an inner degree of freedom (where $m$ and $w$ are continuous parameters). Note that this reparameterization, depending on the inner scaling, has a bias towards sparsity. This has been used by the continuous sparsification methods Spred (Ziyin & Wang, 2023) and PILoT (Jacobs & Burkholz, 2025). In contrast, we do not use the reparameterization for sparsification but for improving the optimization geometry to enable sign alignment and thus seek to counteract its sparsity bias. We propose an inner scaling that promotes sign flips by rebalancing network layers and, correspondingly, the speed in which they learn.
@@
This provably recovers a different sign case than dense training and could thus be seen as complementary (see Figure˜2). Moreover, in combination with a moderate degree of over-parameterization, which tends to be common in PaI, Sign-In fully solves the one-hidden-neuron problem. In line with this insight, we verify in experiments that existing dense-to-sparse methods can be improved by utilizing Sign-In. Most importantly, we demonstrate across multiple experiments that Sign-In significantly enhances the performance of sparse training from scratch, including those using random masks. We attribute this improvement to better sign alignment and the ability to discover flatter minima. However, as we show, Sign-In cannot replicate the same sign flipping mechanism as that of initial dense training, and is still not competitive with the accuracy of dense-to-sparse methods.
@@
Contributions 1) We empirically show that dense-to-sparse training methods identify trainable parameter-mask combinations via early sign alignment allowing them to find flatter minima and achieve state-of-the-art performance. 2) We provide evidence for the following hypothesis: Across dense-to-sparse training masks, the learned signs define an initialization that enables competitive sparse training from scratch. A fully random initialization (that is decoupled from the mask) performs on par or worse than PaI (including random pruning). 3) To learn signs more effectively, we propose Sign-In, a weight reparameterization, which can provably —using Riemannian gradient flow and dynamical systems analysis —recover correct signs in a case that is complementary to overparameterization. 4) Extensive empirical evaluation shows that Sign-In can significantly improve the performance of sparse training from scratch by improving sign alignment and finding flatter minima.
@@ 2 Related Work
Lottery Ticket Hypothesis (LTH) Training sparse neural networks from scratch is a hard problem (Kumar et al., 2024), but Frankle & Carbin (2019) posit that there exist initializations that make sparse networks trainable, i.e., Lottery Tickets. However, their approach to find such initializations by iterative magnitude pruning (IMP) is computationally expensive and its performance does not transfer to larger models, even though lottery tickets should exist in theory da Cunha et al. (2022); Burkholz (2022a, b); Burkholz et al. (2022); Burkholz (2024). Renda et al. (2020) improved its accuracy by Weight Rewinding (WR) and Learning Rate Rewinding (LRR). Both WR and LRR are similarly expensive, but inherit the initialization from denser networks.
@@
Dense-to-sparse methods While dense training has been conjectured to be a critical factor to enter a good loss basin Paul et al. (2023) or quadrant Gadhikar & Burkholz (2024), other methods aim to reduce its duration. Gradual pruning methods like AC/DC (Peste et al., 2021), CAP (Kuznedelev et al., 2024), and WoodFisher (Singh & Alistarh, 2020) prune based on a parameter importance score, interspersed with training. Continuous sparsification methods such as STR (Kusupati et al., 2020), spred (Ziyin & Wang, 2023), and PILoT (Jacobs & Burkholz, 2025) reparameterize a dense neural network to induce sparsity. Although such dense training phases seem to be critical to learn good parameter signs Gadhikar & Burkholz (2024), we propose a complementary method, which also works for pruning at initialization.
An orthogonal class of algorithms, dynamic sparse training, includes RiGL (Evci et al., 2020; Lasby et al., 2023; Chen et al., 2021), SET (Mocanu et al., 2018), MEST (Yuan et al., 2021), CHTs (Zhang et al., 2025) which dynamically modify a sparse mask during training. However, our work focuses on training a fixed mask from scratch.
@@
Pruning at initialization (PaI) A variety of methods have been proposed to sparsify networks at initialization based on different scores that assess parameter importance like saliency in SNIP (Lee et al., 2019) and GraSP (Wang et al., 2020), signal flow in Synflow (Tanaka et al., 2020) and maximal paths (Pham et al., 2023). Other methods include, DPaI (Xiang et al., 2025), Phew(Patil & Dovrolis, 2021), ProsPr (Alizadeh et al., 2022). However, randomly pruned sparse network can perform at par with these methods (Liu et al., 2021; Gadhikar et al., 2023; Ma et al., 2021; Su et al., 2020), calling our ability to identify problem specific masks into question (Frankle et al., 2021; Jain et al., 2024). Kumar et al. (2024) characterize the challenge from an information-theoretic lens, showing that initial dense training, in contrast to PaI, increases effective model capacity. Our goal is not to address the mask identification problem but to improve sparse training given a mask, by a reparameterization.
@@
Reparameterization Reparameterizations have been studied from the perspective of implicit bias and gradient flow (Chizat & Bach, 2020; Li et al., 2022; Woodworth et al., 2020; Gunasekar et al., 2020, 2017; Chou et al., 2024; Vaškevičius et al., 2019; Li et al., 2023; Zhao et al., 2022; Li et al., 2021; Dominé et al., 2024; Jacobs & Burkholz, 2025; Jacobs et al., 2025; Marcotte et al., 2023, 2024; Liang & Montufar, 2025). It explains how overparameterization can improve generalization and how this is affected by various training factors such as the initialization, learning rate, noise, and explicit regularization. An important assumption in these works is that the activation function is linear. This simplification enables the reparameterization to be expressed using mirror flow or Riemannian gradient flow frameworks (Li et al., 2022). In contrast, we employ a reparameterization specifically to induce sign flips in the presence of a non-linear activation function. Our proposed reparameterization has been applied to induce continuous sparsification (Ziyin & Wang, 2023; Jacobs & Burkholz, 2025; Kolb et al., 2025) with the help of explicit regularization. To combat this trend, we rescale the parameters dynamically exploiting a rescale invariance.
@@ 3 Correct Signs Facilitate Sparse Training
Figure 1: Sign learning a $90\%$ sparse ResNet50 on ImageNet. The majority of signs are flipped early for dense-to-sparse methods, upto warmup. Moreover, signs stabilize after warmp-up. In contrast, signs do not stabilize after warmup for sparse training from scratch for different masks.
Paul et al. (2023) have established that iterative pruning methods like Learning Rate Rewinding (LRR) and Weight Rewinding (WR) rely on the effective exploration of the loss landscape during dense overparameterized training. Dense-to-sparse training inherits its performance by training overparameterized in the beginning, identifying a good loss basin. Gadhikar & Burkholz (2024) have shown that this loss basin can be characterized by the weight signs and thus the quadrant of the solution. Our first objective is to show that this finding is not unique to LRR. In fact, it is a common phenomenon that points to a critical limitation of PaI to learn signs.
Our experiments reveal two key findings: (i) Initializing a network with an aligned sign and mask is sufficient for successful sparse training. (ii) Sign alignment happens early during training for dense-to-sparse methods and guides the training trajectory towards flatter minima. Specifically, we study three well-established sparsification methods to draw insights on sign alignment: magnitude based pruning with AC/DC (Peste et al., 2021), continuous sparsification with STR (Kusupati et al., 2020), and dynamic sparse training with RiGL (Evci et al., 2020). The first two methods leverage information from dense training by alternating dense and sparse training and gradual sparsification by soft thresholding respectively. Although, RiGL does not involve a dense training phase, it retains dense gradient information, which we observe serves a similar purpose. For comparison with sparse training, we train a randomly pruned sparse mask from scratch, which has been shown to be competitive with (and even better than) other PaI methods (Liu et al., 2021) and serves as a strong baseline. Noteworthy is the fact that the random mask contains no information about the task. Even though other masks should therefore have an advantage, as we will see later, they cannot shine through because of limited trainability.
@@
Method AC/DC STR RiGL
Baseline $\underline{74.68}$ $\underline{73.65}$ $\underline{73.75}$
Epoch $10$ sign + learned mask $73.97$ $71.7$ $73.32$
Epoch $30$ sign + learned mask $\mathbf{74.89}$ $\mathbf{73.91}$ $73.7$
Final sign + learned mask $74.88$ $73.77$ $\mathbf{73.74}$
Epoch $10$ mag + learned mask $70.93$ $68.01$ $72.30$
Final mag + learned mask $70.94$ $68.35$ $72.40$
Random init + learned mask $70.6$ $68.38$ $71.89$
@@
Signs are learned early To study the alignment between the sparse mask and parameter signs, we track them during dense-to-sparse training of a ResNet50 on ImageNet (see Section˜6 for details). As shown in Figure˜1, the initial signs and final learned signs are significantly different since approximately $50\%$ of the signs are flipped for all methods (blue bar), i.e., the signs at initialization have a close to random overlap with the final learned signs. As training progresses, upto the warmup phase ($10$-th epoch), a significant number of signs flip, as shown by the red bar in Figure˜1. After the warmup phase, fewer signs flip in addition (green bar in Figure˜1). In comparison, training the same masks from scratch starting from a random re-initialization (which is decoupled from the dense-to-sparse training procedure) does not identify stable signs early. The performance of these masks is at par with a random mask trained from scratch (see Table˜8 in the appendix) in the absence of sign alignment. Moreover, the sign flip statistics show a similar pattern as training a random mask from scratch. A significant number of sign flips occur after warmup, and the overlap between initialization, warmup and the final signs is closer to random (see Figure˜1). Across all three dense-to-sparse baselines, we observe that signs are learned early during training and remain largely stable after warmup. In contrast, sparse training does not find stable signs to enable sign alignment, reflecting in its worse performance.
@@
Signs are more informative than magnitude As we verify next, the critical information about the task that is missing from the mask are the parameter signs. To do so, we use signs learned at different stages during training, the learned mask and random magnitudes for each method to initialize a sparse network. As reported in Table˜1, final signs combined with the mask are sufficient for sparse training and achieve performance comparable to the baseline. Moreover, since the signs are identified as early as the $10$-th epoch during training, these early signs (combined with the mask) recover baseline performance with an error of $\approx 1\%$. Full baseline performance can be recovered by using signs from the $30$-th epoch. In contrast, if we initialize the sparse network with magnitudes learned at intermediate epochs, learned mask and random signs, we barely improve over training a randomly intialized sparse mask from scratch Table˜1.
@@
Our results not only suggest that sign alignment is sufficient for sparse training, but it also happens early during dense-to-sparse training. Sparse training from scratch with falls short in comparison.
@@
Flatter loss landscapes with dense-to-sparse methods To dive deeper into the benefits of sign alignment for trainability, we track the maximum eigenvalue of the Hessian of the loss as a proxy of sharpness (Jiang et al., 2019). It can also be interpreted as a proxy measure for the tractability (or conditioning) of the optimization problem. As shown in Figure˜4(b), the dense-to-sparse baselines AC/DC, STR and RiGL have smaller sharpness indicating a flatter loss landscape once signs align during training. In comparison, the random sparse network converges to a sharper minimum.
@@
Our observations lead to the conclusion that dense-to-sparse methods are effective because they inherit critical information from dense training in the form of parameter signs during the early training phase. These signs play a crucial role in dictating the training trajectory of the sparse network. In contrast, this mechanism is absent in training a fixed mask from scratch, resulting in poor performance and a sharper loss landscape during training, which indicates convergence to an inferior loss basin.